OpenStudy (anonymous):

determine where the following two intersect x^2+y^2=13 and y=x^2-1

OpenStudy (anonymous):

y=sprroot of 13-x^2 sprrt13-x^2=x^2-1 13-x^2=x^4-2x^2+1 x^4-x^2-12=0 (x^2-4)(x^2+3)=0 x=2 x=-2 x=0+(i)sqrot3